OpenStudy (anonymous):

How do you find all the zeros of the function: f(x)=x^4-x^3+2x^2-4x-8

OpenStudy (anonymous):

@timo86m

OpenStudy (anonymous):

Factor or graph.

OpenStudy (anonymous):

can you show me how to factor this problem

OpenStudy (anonymous):

factored (x-2)*(x+1)*(x^2+4) ^ notice here you have to take negative sqrt of a the number so it is 2 imaginaries. {x = 2}, {x = -1}, {x = 2*I}, {x = -2*I}

OpenStudy (anonymous):

isn't x^2+4 prime?

OpenStudy (anonymous):

Oh okay http://answers.yahoo.com/question/index?qid=20091115111403AAnQ7Zk I see what you did there :) Gracias!!!!

OpenStudy (anonymous):

De nada :D